OpenStudy (anonymous):

How do i solve a differential equations problem that involves the logistic model equation but the equation is not in the default form?

OpenStudy (anonymous):

for instance, instead is not the normal one: \[\frac{ dP }{ dt } = kP(1-\frac{P}{K})\] but slightly different: \[\frac{ dP }{ dt } = kP(7-\frac{P}{K})\]

OpenStudy (anonymous):

are the k and K constant?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

\[\frac {dP}{P(7 - \frac{P}{k})} = kdt\]

OpenStudy (anonymous):

now solve

OpenStudy (anonymous):

\[\frac{-dP}{7P} + \frac{dP}{7K(\frac{P}{K} - 7)} = -kdt\]